hi thank you for helping me i have 5 problems i do not understand.
5x+2=2x-10
1/2B+4=1/8B+88
8-5P=4P-1
0.2X-8=-2-X
-4-3X=7X-6

Answer by sister_insomnia(8)

5x+2=2x-10
First, subtract 2x from both sides. Which leaves:
3x+2=-10
Then, subtract 2 from both sides:
3x= -12
Finally divide both sides by 3x:
x= -4
-------------------------------------------------
Skip number 2
-------------------------------------------------
8-5P=4P-1
First, add one to each side:
9-5P= 4P
Then, add 5P to both sides:
9=9P
Finally, divide both sides by 9:
1=P
---------------------------------------------
0.2X-8=-2-X
First, add 2 to both sides:
2.2X-8=-X
Then, subtract 2.2X from both sides:
-8=-2.2X
Finally, divide both sides by -2.2:
X=3.63 repeating decimal
-------------------------------------------------
-4-3X=7X-6
First, subtract 7X from both sides:
-4-10X=-6
Then, add -4 to both sides:
-10X=-2
Finally, divide both sides by -10:
X= .2 or 1/5